Use your data from table 8-3 to determine delta H solution for your unknown salt.

Question #4 One way to limit the choices for your unknown is to consider the sign of AHSoln: Considering your results from TaI've tried a few different things but can't seem to come up with the right answers
Answer #1

Trial 1

Heat change in solution (q) = mass of solution × specific heat × temperature change

= (5.418+50) × 4.184× ( 16 - 22)

= 55.418 × 4.184 × (- 6)

= - 1391.2 J.

Heat change in dissolution ( -qsol) = - q = 1391.2 J.

So, \Delta H = (1391/5.148) = 270.24 J/g.

Trial 2

qsol = (5.032+50) × 4.184 × (16 -22)

= 55.032 × 4.184 × (-6)

= - 1381.52 J.

So, Heat change in dissolution (qsol) = - q = 1381.52 J.

\Delta Hsol = ( qsol/m) = (1381.52/5.032) = 274.54 J/g.